1) Find the chi-square critical value for a right-tailed hypothesis test about variance at the 1% significance level with a sample size of 15
Answer #1

We have given,

Sample size =n=15

Degree of freedom =n-1=15-1=14

Chi square critical value =29.141............by using chi square critical value table or Excel =CHIINV(0.01,14)